# Tilus: A Virtual Machine for Arbitrary Low-Precision GPGPU Computation in LLM Serving

### v2

@@ 3.2. An Example of Virtual Machine Program
Figure 2 illustrates a low-precision matrix multiplication in the virtual machine. Matrix multiplication is defined as $C_{M,N}=A_{M,K}\times B_{K,N}$, where $A$ and $B$ are of type float16 (a 16-bit floating-point number (IEEE Standards Association, 2019)) and int6 (a 6-bit signed integer), respectively. The kernel performs matrix multiplication with given M, N, and K, where each thread block computes a BM $\times$ BN tile of the C matrix (Line 1). Therefore, a grid of (M / BM, N / BN) thread blocks must be launched (Line 2). Inside the kernel, the BlockIndices instruction retrieves the thread block indices bi and bj (Line 3), which determine the offset (bi * BM, bj * BN) for computing the corresponding C tile. Three tensor views are created for the input and output tensors in global memory by specifying their addresses and shapes (Line 4-6). Then, a register tensor of type f16[16, 8] is created with the following layout:
@@
It distributes $16\times 8=128$ elements across 32 threads, with each thread storing 4 elements (Line 7). This layout is composed of three primitive layouts (Section 4) and aligns with the C matrix layout used by the mma.m16n8k16 tensor core instruction in PTX (NVIDIA Corporation, 2024b). The reduction loop over the k dimension (Line 8-13) repeatedly loads tiles of A and B from global memory into registers and performs matrix-multiply accumulate (mma). For each iteration, we first load a f16[16, 16] tile from global memory to register with a LoadGlobal instruction (Line 9). The layout of the loaded register tile is specified and required by the tensor core instruction. The offset parameter specifies the position of the loaded tile within the global tensor. Loading tensor B, which has data type int6, involves a more complex process, detailed in Section 6. We summarize the high-level ideas here. As a pre-processing step before launching the kernel, the weight tensor’s layout in global memory is transformed from i6[K, N] to u8[K / BK, N / BN, BK * BN * 6 / 8], enabling efficient loading via the LoadGlobal instruction (the ‘Change Layout’ step in Figure 2 (b)). Next, in the kernel, the transformed tile is loaded into a register tensor (Line 10) and then reinterpreted to a tensor with a different data type and layout (Line 11). This reinterpretation is valid because both tensors are stored across the same number of threads (32), with each thread holding exactly 24 bits (i.e., 3 $\times$ u8 or 4 $\times$ i6), as shown in Figure 2 (c). Following this, the i6 tensor is cast to an f16 tensor (Line 12), which is then fed to the tensor core to perform matrix-multiply accumulate (mma) (Line 13). Finally, the accumulation tensor is cast from f32 to f16 and stored in global memory (Line 14-15). For simplicity, this program does not utilize shared memory and omits optimizations such as software pipelining (Huang et al., 2023); additionally, each k-iteration performs only a single tensor core instruction (NVIDIA Corporation, 2024a). An optimized implementation can be found in Appendix B.
The following sections introduce the three core components of the proposed VM. Section 4 introduces an algebraic layout formulation to systematically define how the elements of a tile are stored in the registers among the block threads. Section 5 introduces the thread-block-level programming model with a hierarchical memory space exposed explicitly. Section 6 introduces the native support for arbitrary low-precision data types to address the growing demand for low-precision computation in deep learning workloads.
4. Algebriac Layout System
@@
The virtual machine exposes a hierarchical memory space with global memory, shared memory, and registers to programmers. We need a way to model the mapping between the logical index of a tensor element and the location of the corresponding element in memory for all three memory scopes. Such a mapping is usually called the layout of the tensor. Figure 3 illustrates an example of the layout used by a tensor core instruction: mma.m16n8k8.f32.f16.f16.f32 D, A, B, C. It performs the following computation: $D_{16,8}=A_{16,8}B_{8,8}+C_{16,8}$ where $A,B,C,D$ are tensors stored in thread registers and distributed across the $32$ threads in a warp. Since the elements are spread across different threads, we refer to this layout as a distributed layout (Tillet et al., 2019). Such a layout can be defined as a function $f$ that maps a thread index $t$ and a local index $i$ within that thread to the logical index $f(t,i)$ of the corresponding tensor element. For example, the layout in Figure 3 can be represented as:
@@
Here, $t$ ranges from 00 to $31$, and $i$ ranges from 00 to $3$. The function $f(t,i)$ represents the logical index of element $i$ in thread $t$. Since all threads in a thread block have access to shared and global memory, this formalization can also be used to describe their layouts under a single-thread assumption. That is, by setting $t=0$, we define $f(0,i)$ as the logical index of the element at address $i$ in shared or global memory.
@@ 4.1. Parameterized Primitive Layouts
With the formal definition of layout, we introduce parameterized primitive layouts that serve as the fundamental building blocks of our layout algebra. Given a tile [^2] We use the terms tile and tensor interchangeably. with shape $(n_{1},n_{2})$, there are two primary ways to store it: (1) storing all $n_{1}n_{2}$ elements in a single thread, or (2) distributing all elements across $n_{1}n_{2}$ threads, where each thread holds a single element. We refer to the first type as local layouts, denoted as local(n1, n2), and the second as spatial layouts, denoted as spatial(n1, n2). This concept naturally extends to tiles with arbitrary dimensions. Figure 4 illustrates these two primitive layouts. The local(2, 3) layout maps the $i$-th local element of thread $t$ to the logical index $(i/3,i\ \%\ 3)$, while the spatial(2, 3) layout maps it to $(t/3,t\ \%\ 3)$. We observe that all complex layouts used in modern deep learning workloads and GPUs can be constructed using these two primitive layouts. In the next section, we introduce methods for composing these layouts to define more complex layouts.
4.2. Layout Composition
Figure 5. Complex layouts can be constructed through composition. In the figure, layout (c) is composed of layouts (a) and (b), while layout (e) is composed of layouts (d) and (a).
@@ 4.2. Layout Composition
The layouts used in modern deep learning workloads, as well as those defined by hardware instructions, typically exhibit a hierarchical structure. Consider layout (c) in Figure 5 as an example. This layout has shape $(4,6)$, storing $24$ elements across $6$ threads, with each thread holding four elements. We denote the four elements stored in each thread as $a_{0},a_{1},a_{2},a_{3}$. Comparing its first two rows with the last two, we observe a similar structure, except that the last two rows store elements in $a_{2}$ and $a_{3}$ instead of $a_{0}$ and $a_{1}$. To model this structural invariance, layout (c) can be viewed as a composition of layouts (a) and (b), where each element in layout (a) represents a tile with layout (b). In fact, layouts (a) and (b) can be combined to represent layout (c) as follows:
@@
where $60\leq t<$ 6, $40\leq i<$ 4, $\odot$ denotes the element-wise product, and $(2,6)$ represents the shape of layout (c). This composition principle can be generalized. Given two layouts $f$ and $g$ with the same number of dimensions, we define their composition $h=f\circ g$ as
@@
where $T_{g},N_{g},S_{g}$ represent the number of threads, the number of local elements per thread, and the shape of layout $g$, respectively. We can prove that composition is associative, meaning that for any three layouts $f$, $g$, and $h$, the equality $f\circ(g\circ h)=(f\circ g)\circ h$ holds. However, composition is not commutative, meaning that, in general, $f\circ g\neq g\circ f$. Spatial and local layouts follow a row-major ordering for threads and local elements, respectively. Using composition, we can construct their column-major counterparts, column_spatial(...) and column_repeat(...), as demonstrated by layout (e) in Figure 5. Returning to the tensor core instruction layout in Figure 3, it can be expressed as a composed layout local(2, 1).spatial(8, 4).local(1, 2). Using layout composition, we can also define its inverse operation. If $h=f\circ g$, we define $f=h/g$ as the result of dividing layout $h$ by layout $g$. For example, dividing local(2, 4) by local(1, 2) results in local(2, 2). In Appendix A, we formally define layouts and prove that under the composition operator, they form a semigroup with an identity element, making them a monoid.
@@ 5. Thread-Block-Level Programming Model
Modern GPU programming models, such as PTX (NVIDIA Corporation, 2024b) and CUDA (NVIDIA Corporation, 2024a), define operations at the thread level, following the Single-Instruction-Multiple-Thread (SIMT) paradigm (Aamodt et al., 2018). To simplify GPU programming, we adopt the thread-block-level programming model, defining operations at the granularity of thread blocks rather than individual threads. Additionally, building on the layout system introduced previously, we propose explicitly exposing the hierarchical memory structure in modern GPUs, enabling fine-grained memory control while reducing programming complexity. We refer to this model as Single-Instruction-Multiple-Block (SIMB).
@@ 5.1. State Space and Type System
The virtual machine supports three types of variables. Scalar variables store individual values, such as integers (e.g., int32) or floating-point numbers (e.g., float16). Pointer variables store memory addresses rather than direct data values. Tensor variables represent multi-dimensional arrays, with types that specify their shape, element type, memory scope, and layout. Tensors reside in different memory scopes, including global memory, shared memory, and registers. The tensor layout determines how high-dimensional tensor elements are mapped to linear memory. All variables in the virtual machine operate at the thread-block level, meaning that all threads within a block collaboratively maintain them.
@@ 5.2. Program Structure and Control Flow
Figure 6. Virtual machine program structure. A virtual machine program contains parameters and a body. The body is a list of control-flow statements or block-level instructions.
@@
Table 1. The thread-block-level instruction set of the virtual machine. Each instruction specifies an operation applied to the entire thread block. Parameters enclosed in [...] are optional. Instructions that return a new register tensor also have an in-place variant, which writes the result to an existing register tensor using the out parameter instead of creating a new tensor.
Figure 6 illustrates the structure of a VM program. Each program consists of a program name, a grid shape, a list of parameters, and a program body. The grid shape is specified as a list of expressions enclosed in <...>, where each expression is either a positive integer or an integer expression based on the program parameters. If the grid shape contains parameter-based expressions, its dimensions are determined at runtime based on the program’s launch arguments. The program body consists of a sequence of statements, including if-else statements, range-based for-loops, and while-loops. Unlike other low-level virtual machines (NVIDIA Corporation, 2024b) or instruction set architectures (ISAs) (NVIDIA Corporation, 2024c), our virtual machine does not abstract control-flow statements into jump instructions. Instead, it retains high-level control structures to improve readability and ease of programming for human developers. In addition to control-flow statements, individual instructions can also serve as statements. Most functionalities provided by the virtual machine are implemented as instructions in the thread-block-level instruction set.
@@ 5.3. Thread-Block-Level Instruction Set
Each instruction in the VM’s instruction set operates at the thread-block level rather than the thread level. Table 1 shows a list of the instructions in the instruction set, with the signature of each instruction and a brief description of the instruction semantics. These instructions allocate tensors with specific data types, shapes, and layouts in designated memory spaces (e.g., global memory, shared memory, registers), transfer tensors between memory spaces, and perform computations or transformations on register tensors. Modern processors execute instructions out of order (Tomasulo, 1967), meaning that a subsequent instruction may begin execution before the current one completes, provided there are no dependencies between them. Similarly, the execution of instructions in our virtual machine exhibits this behavior: certain subsequent instructions may begin execution before the current instruction completes. Generally, this behavior does not pose significant issues. However, an exception occurs when two instructions access the same region of shared or global memory, and the second instruction depends on the completion of the first. In such cases, a Synchronize instruction must be inserted to ensure all preceding instructions complete before subsequent ones execute.
@@ 6. Arbitrary Low-Precision Data Types
Modern processors use bytes (8 bits each) as the smallest processing unit. As a result, standard data types in modern programming languages typically have bit widths that are multiples of 8. However, the high computational and memory demands of LLMs make low-precision data types (less than 8 bits) essential for reducing resource consumption. This section describes how the virtual machine efficiently supports low-precision data types.
@@ 6.1. Storage of Low-Precision Data
Since modern processors, including CPUs and GPUs, use bytes as the smallest unit for memory access and computation, we store low-precision data (fewer than 8 bits per element) compactly within bytes, as shown in Figure 7. Compact storage eliminates bit gaps between consecutive low-precision values, which may result in a single value spanning two uint8 entries (e.g., b[1] in Figure 7). Bitwise operations are employed to extract, manipulate, and store low-precision values within packed byte arrays. To load a low-precision value, we first extract relevant bits using bitwise AND, adjust their position with bitwise SHIFT operations, and finally combine separated parts using bitwise OR if the value spans multiple bytes. Similarly, to store a low-precision value, we first clear the target bit positions using a bitwise mask, then insert the new value using bitwise OR while preserving the other bits. Low-precision data is cast to standard data types before arithmetic computations and cast back afterward. While these methods enable support for arbitrary bit-width data types, they are often inefficient and serve only as a fallback mechanism. More efficient handling of low-precision data is necessary for LLM serving.

### v3

@@ 3.2. An Example of Tilus Program
Figure 2 illustrates a low-precision matrix multiplication in Tilus. Matrix multiplication is defined as $C_{M,N}=A_{M,K}\times B_{K,N}$, where $A$ and $B$ are float16 (a 16-bit floating-point number (IEEE Standards Association, 2019)) and int6 (a 6-bit signed integer), respectively. The kernel performs matrix multiplication with dimensions M, N, and K, where each thread block computes a BM $\times$ BN tile of the C matrix (Line 1). Therefore, a grid of (M / BM, N / BN) thread blocks must be launched (Line 2). Inside the kernel, the BlockIndices instruction retrieves the thread block indices bi and bj (Line 3), which determine the offset (bi * BM, bj * BN) for computing the corresponding C tile. Three tensor views are created for the input and output tensors in global memory by specifying their addresses and shapes (Line 4-6). Then, a register tensor of type f16[16, 8] is created with the following layout:
@@
It distributes $16\times 8=128$ elements across 32 threads. Each thread stores 4 elements (Line 7). This layout is composed of three primitive layouts (Section 4) and aligns with the C matrix layout used by the mma.m16n8k16 tensor core instruction in PTX (NVIDIA Corporation, 2024b). The reduction loop over the k dimension (Line 8-13) repeatedly loads tiles of A and B from global memory into registers and accumulates their product. For each iteration, we first load a f16[16, 16] tile from global memory to register with a LoadGlobal instruction (Line 9). The layout of the loaded register tile is specified and required by the tensor core instruction. The offset parameter specifies the position of the loaded tile within the global tensor. Loading tensor B, with data type int6, involves a more complex process, detailed in Section 7. We summarize the high-level ideas here. As a pre-processing step before launching the kernel, the weight tensor’s layout in global memory is transformed from i6[K, N] to u8[K / BK, N / BN, BK * BN * 6 / 8], enabling efficient loading via the LoadGlobal instruction (the ‘Change Layout’ step in Figure 2 (b)). Next, in the kernel, the transformed tile is loaded into a register tensor (Line 10) and then reinterpreted to a tensor with a different data type and layout (Line 11). This reinterpretation is valid because both tensors are stored across the same number of threads (32), with each thread holding exactly 24 bits: 3 $\times$ u8 or 4 $\times$ i6, as shown in Figure 2 (c). Following this, the i6 tensor is cast to an f16 tensor (Line 12), which is then fed to the tensor core to perform matrix-multiply accumulate (mma) (Line 13). Finally, the accumulation tensor is cast from f32 to f16 and stored in global memory (Line 14-15). For simplicity, this program does not use shared memory and omits optimizations like software pipelining (Huang et al., 2023). Additionally, each k-iteration performs only a single tensor core instruction (NVIDIA Corporation, 2024a).
The following sections introduce the three core components of Tilus. Section 4 introduces an algebraic layout formulation to systematically define how the elements of a tile are stored in the registers among the block threads. Section 6 introduces the thread-block-level programming model with a hierarchical memory space exposed explicitly. Section 7 introduces the native support for arbitrary low-precision data types to address the growing demand for low-precision computation in deep learning workloads.
4. Algebraic Layout System
@@
Tilus exposes a hierarchical memory space to programmers, comprising global memory, shared memory, and registers. We need a way to model the mapping between the logical index of a tensor element and the location of the corresponding element in memory for all three memory scopes. Such a mapping is usually called the layout of the tensor. Of the memory scopes, the layout for register tensors is the most complicated. Figure 3 illustrates an example of the layout used by a tensor core instruction: mma.m16n8k8.f32.f16.f16.f32 D, A, B, C. It performs the following computation: $D_{16,8}=A_{16,8}B_{8,8}+C_{16,8}$ where $A,B,C,D$ are tensors stored in thread registers and distributed across the $32$ threads in a warp. Since the elements are spread across different threads, we refer to this layout as a distributed layout (Tillet et al., 2019). Such a layout can be defined as a function $f$ that maps a thread index $t$ and a local index $i$ to the logical index $f(t,i)$ of the tensor element. For example, the layout in Figure 3 can be represented as:
@@
Here, $t$ ranges from 0 to $31$, and $i$ from 0 to $3$. The function $f(t,i)$ represents the logical index of the element stored in the local element $i$ in thread $t$.
@@ 4.1. Parameterized Primitive Layouts
With the formal definition of layout, we introduce parameterized primitive layouts that serve as the fundamental building blocks of our layout algebra. Given a tile [^2] We use the terms tile and tensor interchangeably. with shape $(n_{1},n_{2})$, there are two primary ways to store it: (1) to store all $n_{1}n_{2}$ elements in a single thread, or (2) to distribute all elements across $n_{1}n_{2}$ threads, with each thread holding a single element. We refer to the first type as local layouts, denoted as local(n1, n2), and the second as spatial layouts, denoted as spatial(n1, n2). This concept naturally extends to tiles with arbitrary dimensions. Figure 4 illustrates these two primitive layouts. The local(2, 3) layout maps the $i$-th local element of thread $t$ to the logical index (i/3,i% 3)(i/3,i\ $\%\$3), while the spatial(2, 3) layout maps it to (t/3,t% 3)(t/3,t\ $\%\$3). We observe that the layouts for all common operators in LLMs can be constructed using these two primitive layouts. In the next section, we introduce the Kronecker product of layouts to construct more complex layouts.
4.2. Kronecker Product
Figure 5. Examples of Kronecker products over layouts. In the figure, layout (c) is the product of layouts (a) and (b), while layout (e) is the product of layouts (d) and (a).
@@ 4.2. Kronecker Product
The layouts used in modern deep learning workloads, as well as those defined by hardware instructions, typically exhibit a hierarchical structure. Consider layout (c) in Figure 5 as an example. This layout has shape $(4,6)$, storing $24$ elements across $6$ threads. Each thread holds four elements. We denote the four elements stored in each thread as $a_{0},a_{1},a_{2},a_{3}$. Comparing its first two rows with the last two, we observe a similar structure, except that the last two rows store elements in $a_{2}$ and $a_{3}$ instead of $a_{0}$ and $a_{1}$. To model this structural invariance, layout (c) can be viewed as a Kronecker product of layouts (a) and (b), with each element in layout (a) representing a tile with layout (b). Indeed, layouts (a) and (b) can be multiplied to represent layout (c):
@@
where $0\leq t<6$, $0\leq i<4$, $\odot$ denotes the element-wise product, and $(2,6)$ represents the shape of layout (c). The Kronecker product can be generalized. Given two layouts $f$ and $g$ with the same number of dimensions, we define their Kronecker product, $h=f\otimes g$ as
@@
where $T_{g},N_{g},S_{g}$ represent the number of threads, the number of local elements per thread, and the shape of layout $g$, respectively. We can prove that this operation is associative, meaning that for any three layouts $f$, $g$, and $h$, the equality $f\otimes(g\otimes h)=(f\otimes g)\otimes h$ holds. However, the operation is not commutative, meaning that, in general, $f\otimes g\neq g\otimes f$. Spatial and local layouts follow a row-major ordering for threads and local elements, respectively. Using this operation, we can construct their column-major counterparts, column_spatial(...) and column_local(...), as demonstrated by layout (e) in Figure 5. Returning to the tensor core instruction layout in Figure 3, it can be expressed as a Kronecker product local(2, 1).spatial(8, 4).local(1, 2). We can also define its inverse operation. If $h=f\otimes g$, we define $f=h/g$ as the result of dividing layout $h$ by layout $g$. For example, dividing local(2, 4) by local(1, 2) results in local(2, 2). We also refer to the Kronecker product as layout composition in some contexts.
5. Unified Layout Representation
Figure 6. Example of the unified layout representation.
[…]
We use a unified representation for all layouts of register tensors in Tilus. This representation gives each layout four attributes: shape, mode_shape, spatial_modes, and local_modes. The shape is a sequence of integers that defines the shape of the register tensor. We can split each dimension of the register tensor into some sub-dimensions (we call them mode following prior work (Hagedorn et al., 2023)) and concatenate these sub-dimensions to get the mode_shape. Then we use spatial_modes and local_modes to specify the sub-dimensions assigned to spatial threads and to the local storage of each thread. The dimension split-distribute-merge method uniquely defines a register layout.
Figure 6 shows an example of a layout and how we map the logical index of a register tensor element to the pair of thread_index and local_index. There are three steps: 1) split dimensions, 2) distribute sub-dimensions; and 3) merge sub-dimensions. Given a logical index [i, j], we first 1 split each index into the indices of its sub-dimensions (i.e., [i0, i1, i2] for i and [j0, j1, j2] for j) with the unravel operation. After that, we 2 distribute the sub-dimension indices to get the indices for spatial threads (i.e., [i2, j1]) and local storage ([i0, j0, i1, j2]). Finally, we 3 convert the multi-dimensional indices for threads and local storage into linear index to get the thread index and local index. The ravel and unravel functions are used to convert between multi-dimensional index in a grid with given shape and its row-major linear index. For example: unravel(i, [4, 2, 8]) = [i / 16, i / 8 % 2, i % 8], and ravel([i2, j1], [8, 4]) = i2 * 4 + j1.
The layouts represented in this form are closed under the Kronecker product, meaning that the product result of two layouts in this form can also be represented in this form.
@@ 6. Thread-Block-Level Programming Model
Modern GPU programming models, such as PTX (NVIDIA Corporation, 2024b) and CUDA (NVIDIA Corporation, 2024a), define operations at the thread level, following the Single-Instruction-Multiple-Thread (SIMT) paradigm (Aamodt et al., 2018). To simplify GPU programming, we adopt the thread-block-level programming model, defining operations at the granularity of thread blocks rather than individual threads. Additionally, building on the layout system introduced previously, we propose explicitly exposing the hierarchical memory structure in modern GPUs, enabling fine-grained memory control while reducing programming complexity. We refer to this model as Single-Instruction-Multiple-Block (SIMB). In this section, we will introduce the type system, program structure, and instruction set of Tilus.
@@ 6.1. State Space and Type System
Tilus supports three variable types. Scalar variables store individual values, such as integers (e.g., int32) or floating-point numbers (e.g., float16). Pointer variables store memory addresses rather than direct data values. Tensor variables represent multi-dimensional arrays, with types that specify their shape, element type, memory scope, and layout. Tensors reside in different memory scopes, including global memory, shared memory, and registers. The tensor layout determines how high-dimensional tensor elements are mapped to linear memory. All variables in Tilus operate at the thread-block level, meaning that all threads within a block collaboratively allocate and maintain their state.
@@ 6.2. Program Structure and Control Flow
Figure 7. A Tilus program contains parameters and a body. The body is a list of control-flow statements or block-level instructions. The majority of functionality, such as tensor allocation and computation, is provided by instructions.
@@
Table 1. The thread-block-level instruction set of Tilus’s virtual machine. Each instruction specifies an operation applied to the entire thread block. Parameters enclosed in [...] are optional. Instructions that return a new register tensor also have an in-place variant, which writes the result to an existing register tensor using the out parameter instead of creating a new tensor.
Figure 7 illustrates the structure of a Tilus program. Each program consists of a program name, a grid shape, a list of parameters, and a program body. The grid shape is specified as a list of expressions enclosed in <...>, where each expression is either a positive integer or an integer expression based on the program parameters. If the grid shape contains parameter-based expressions, its dimensions are determined at runtime based on the program’s launch arguments. The program body consists of a sequence of statements, including if-else statements, range-based for-loops, and while-loops. Unlike other low-level virtual machines (NVIDIA Corporation, 2024b) or instruction set architectures (ISAs) (NVIDIA Corporation, 2024c), our virtual machine does not abstract control-flow statements into jump instructions. Instead, it retains high-level control structures to improve readability and ease of programming for human developers. In addition to control-flow statements, individual instructions can also serve as statements. Most of Tilus’s functionality is implemented as instructions within its instruction set.
@@ 6.3. Thread-Block-Level Instruction Set
Each instruction in the Tilus’s instruction set operates at the thread-block level rather than the thread level. Table 1 shows a list of the instructions in the instruction set, with the signature of each instruction and a brief description of the instruction semantics. These instructions allocate tensors with specific data types, shapes, and layouts in designated memory spaces (e.g., global memory, shared memory, registers), transfer tensors between memory spaces, and perform computations or transformations on register tensors. The execution model of modern GPUs allows different warps to execute different instructions at the same time. Similarly, the execution of instructions in Tilus exhibits this behavior: certain subsequent instructions may begin execution before the current instruction completes, resulting in multiple block-level instructions being in-flight simultaneously. Generally, this behavior does not pose significant issues. However, an exception occurs when two instructions access the same region of shared or global memory, and the second instruction depends on the completion of the first. In such cases, a Synchronize instruction must be inserted to ensure all preceding instructions complete before subsequent ones execute. Instructions like Print are used for debugging.
@@ 7. Arbitrary Low-Precision Data Types
Modern processors use bytes (8 bits each) as the smallest processing unit. As a result, standard data types in modern programming languages typically have bit widths that are multiples of 8. However, the high computational and memory demands of LLMs make low-precision data types with less than 8 bits essential for reducing resource consumption. This section describes how Tilus efficiently supports low-precision data types with bit width from 1 to 8.
@@ 7.1. Storage of Low-Precision Data
[…]
Since modern processors, including CPUs and GPUs, use bytes as the smallest unit for memory access and computation, we store low-precision data (fewer than 8 bits per element) compactly within bytes, as shown in Figure 8. Compact storage eliminates bit gaps between consecutive low-precision values, which may result in a single value spanning two uint8 entries (e.g., b[1] in Figure 8). Bitwise operations are employed to extract, manipulate, and store low-precision values within packed byte arrays. To load a low-precision value, we first extract relevant bits using bitwise AND, adjust their position with bitwise SHIFT operations, and finally combine separated parts using bitwise OR if the value spans multiple bytes. Similarly, to store a low-precision value, we first clear the target bit positions using a bitwise mask, then insert the new value using bitwise OR while preserving the other bits. Low-precision data is cast to standard data types before arithmetic computations and is cast back afterward. While these methods enable support for arbitrary bit-width data types, they are often inefficient. They serve only as a fallback mechanism. More efficient handling of low-precision data is necessary for LLM serving.